# ARCTraj: A Dataset and Benchmark of Human Reasoning Trajectories for Abstract Problem Solving

### v2

@@ 1. Introduction
Understanding and modeling how humans reason and solve problems, rather than reproducing only their final answers, remains a central challenge in AI (collins2024building). Such problem-solving involves conceptual abstraction, attention shifts, and the use of flexible strategies, yet these abilities remain difficult for machines to emulate (reisberg2013oxford; ho2022people). The Abstraction and Reasoning Corpus (ARC) (chollet2019ARC) was introduced to benchmark these capabilities through grid-based tasks where solvers must infer rules from a small set of input-output examples.
@@
While the benchmark itself is well designed, most ARC research has focused on reproducing outputs rather than modeling reasoning. Early program synthesis approaches (banburski2020dreaming; xu2023graphs; ouellette2024towards; rocha2025ilpar; ferre2025madil) struggled to infer solution strategies because the benchmarks provide limited guidance for program construction. To address this, several studies have attempted to extract auxiliary cues from input grids or from intermediate program execution results; however, these approaches have remained task-specific and brittle. More recently, LLM-based methods (tan2024llms; lei2025reasoning; pourcel2025soar) have leveraged prior knowledge to generate diverse candidate programs, showing improved flexibility but still facing the fundamental challenge of an enormous search space without explicit reasoning guidance.
@@
In parallel, test-time training approaches (akyurek2025tttarc; franzen2024architect; li2025combining) aim to enable on-the-fly adaptation during inference; however, their learning signals still rely on static supervision rather than the evolving structure of reasoning. As a result, existing ARC approaches (e.g., program synthesis, neuro-symbolic reasoning, and test-time learning) remain limited by their reliance on static supervision and their inability to capture the unfolding of human reasoning over time.
@@
To address this limitation, we present ARCTraj, a large-scale dataset of human reasoning trajectories collected while solving ARC tasks. Each trajectory captures a temporally ordered sequence of object-level actions (i.e., moving, rotating, and flipping objects) that transform an input grid into its correct output. By capturing these multi-step trajectories, ARCTraj provides explicit reasoning supervision that connects human perception, abstraction, and strategy formation, offering a structured alternative to unconstrained program search. These logs were collected via the O2ARC web interface (shim2024o2arc), which is designed to support natural human interaction with ARC problems. Each trajectory is annotated with metadata, including timestamps, task identifiers, and success labels, enabling both learning and analysis.
@@
Figure 1. Overview of the ARCTraj data collection process. Users solve ARC tasks through the O2ARC platform by interacting with grid-based objects. Their actions are recorded step-by-step to form semantically rich, temporally ordered trajectories.:
@@
Compared to existing human ARC datasets, ARCTraj offers several advantages. Whereas H-ARC (legris2024harc) captures pixel-level edit logs and the ARC-Interactive-History-Dataset (neoneye2024arc-interactive) records low-level cell and operation sequences, ARCTraj provides object-centric actions with consistent formatting and semantic structure across all 400 ARC-AGI-1 training tasks. This object-level abstraction reflects how humans group meaningful visual units and apply conceptual transformations rather than pixel-level edits. Such abstraction preserves strategic intent and eliminates redundant operations, resulting in trajectories that more accurately represent human reasoning.
@@
Ultimately, ARCTraj shifts the focus of ARC research from reproducing outputs to modeling reasoning itself. By learning from human trajectories, models can acquire structured strategies that reflect System 2 thinking (kim2024system) and generalize across domains, such as program synthesis, robotics, and data transformation. We expect ARCTraj to promote research on transferable reasoning strategies beyond training tasks.
@@ 2. Related Work
ARC (chollet2019ARC), also known as ARC-AGI, is a benchmark designed to test human-like generalization in abstract reasoning tasks. Each task consists of a few input–output grid pairs, requiring solvers to induce and apply conceptual transformations from limited examples. It has inspired research across multiple paradigms, including program synthesis (barke2024hysynth; bober2024neural; butt2024codeit), neuro-symbolic reasoning (batorski2025nsa; liao2024compressarc; xu2023graphs), and test-time training (akyurek2025tttarc; franzen2024architect; li2025combining), many of which were featured in the ARC Prize 2024 Technical Report (chollet2024arcprize). However, ARC remains static; it provides only input–output pairs, offering no insight into the intermediate reasoning that drives human problem-solving. As a result, most studies evaluate solution generalization rather than the reasoning process itself. Understanding this process is crucial for modeling human-like abstraction. In this work, we use the term ARC to refer to the 400 training tasks from the ARC-AGI-1 benchmark, which form the foundation for all trajectories in ARCTraj.
@@
Several efforts have captured traces of human reasoning in ARC tasks. LARC (acquaviva2022communicating) collects natural language explanations, providing semantic insight but lacking action-level detail. Fast and Flexible (johnson2021fast) and H-ARC (legris2024harc) record pixel-level edits, while the ARC-Interactive-History-Dataset (neoneye2024arc-interactive) logs cell-level actions through the BrainGridGame interface. Although these datasets represent progress, their low-level granularity and inconsistent coverage make it challenging to extract structured reasoning or integrate data into learning frameworks. They reveal what humans edit, but not how strategic reasoning evolves. A higher-level, structured view is still missing. These gaps motivate ARCTraj, which captures object-level, temporally structured trajectories aligned across all ARC tasks to support both analysis and model training.
Figure 2. Example of a single trajectory log in ARCTraj. Each action in the “actionSequence” specifies its category and operation, along with the associated grid and object state, forming a structured state–action unit. :
@@ 3.1. Data Collection and Action Structure
Fig. 2 illustrates the trajectory structure in ARCTraj and the components of each action. Each log represents one human solving attempt and contains an ordered sequence of symbolic actions (actionSequence) recorded through the O2ARC interface (shim2024o2arc). Every action is represented as a triplet $\langle\text{category},\text{object},\text{operation}\rangle$ associated with its grid state and timestamp.
@@
These elements define an MDP-compatible format $\langle s_{t},a_{t},s_{t+1}\rangle$, where $s_{t}$ encodes the grid configuration and $a_{t}$ represents the above triplet (lee2024arcle). Formally, the reward $r_{t}\in\{1,0\}$ indicates whether the current grid $s_{t}$ matches the correct output. By preserving order and object identity, ARCTraj captures evolving reasoning processes beyond static outcomes, enabling analysis of strategies such as hypothesis testing, selective attention, and object-level abstraction.
@@ 3.2. Comparative Dataset Statistics
Among existing human ARC datasets, the most representative one is H-ARC (legris2024harc), which also records human solving processes on ARC tasks. However, the two datasets differ notably in both task coverage and the representation of actions. H-ARC collected trajectories from both the training and evaluation splits of ARC-AGI-1 and primarily logs pixel-level edits, where each action merges pixel selection and color change.
@@ 4.1. Formalizing ARC Solvers with ARCTraj
Figure 3. Overview of how ARCTraj analyses inform ARC solving. The ARC solver $f_{\theta}$ predicts $\hat{y}^{\text{test}}$ given $x^{\text{test}}$, demonstration examples $\mathcal{D}_{\text{demo}}$, and auxiliary knowledge $\mathcal{A}$ derived from human trajectories. ARCTraj provides three structured components of $\mathcal{A}$, (1) selection biases, (2) color origins, and (3) shared intentions, that capture different stages of human reasoning and can be integrated into model learning. :
@@ 5. Auxiliary Knowledge from ARCTraj
To structure this investigation, we frame three research questions (RQs) that correspond to the main stages of human problem solving, from perceptual attention to hypothesis formation and strategic abstraction (reisberg2013oxford; ho2022people). Each RQ highlights a distinct aspect of reasoning behavior, together forming a coherent picture of how humans decompose, transform, and recombine patterns in ARC task solving, providing insights that bridge human cognitive processes and AI reasoning frameworks (collins2024building).
@@ 5.1. Biases in Human Grid Selections
Refer to caption:
Refer to caption:
Figure 4. Distributions of human selection behavior in ARC tasks. Left: Selections are concentrated in compact shapes ranging from $1\times 1$ to $3\times 3$, with square- and bar-shaped regions dominating. Right: Most selections cover fewer than 20 pixels, reflecting a preference for local and perceptually salient regions. :
@@ 5.3. Shared Intentions and Strategy Patterns
Figure 5. Uniqueness analysis of human reasoning trajectories. According to the left panel, most ARC tasks show low unique trajectory ratios, indicating that human solvers often converge on similar reasoning paths. The right panel shows a representative low-uniqueness task (Task c0f76784; see Appendix A), where overlapping solution routes appear in the state-space graph. :
@@
Figure 6. Preprocessing ARCTraj for downstream learning. For RL environments such as ARCLE, ARCTraj is filtered to retain only operation actions and is mapped to a Markovian state-action format using grid, object, and operation. For sequence models such as Decision Transformer, only grid and operation are used, omitting intermediate objects and environment interaction.:
@@ 6.1. Current ARC Solvers and Key Trends
Recent research has explored diverse approaches to solving ARC tasks by framing visual reasoning and pattern transformation as learning or generation problems. As shown in Fig. 6, ARCTraj data are reformatted for downstream learning across solver paradigms. For reinforcement learning environments such as ARCLE (lee2024arcle), trajectories are mapped to Markovian state–action pairs using grid, object, and operation. For sequence models such as Decision Transformer (park2023unraveling), only grid and operation are retained to simplify intermediate interactions while preserving reasoning context. This unified preprocessing enables direct comparison of solvers trained on shared human reasoning traces.
@@
PPO (lee2024arcle) Reinforcement Learning Solve 55–70% Demonstrated online training in ARC-like MDPs.
World Model (lee2024analogical) Reinforcement Learning Solve 38–100% Enabled analogical generalization via latent dynamics.
Decision Transformer (park2023unraveling; kim2025intention) Sequential Modeling Solve 59–90% Learned trajectory-conditioned policies with inferred intentions.
Diffusion (kim2024diffusion) Generative Modeling Solve 77–92% Generated intermediate states for plan synthesis.
GFlowNet (hwang2025gfn) Generative Modeling Augment 10–100% Sampled diverse goal-directed solution trajectories.
Reinforcement learning methods (lee2024arcle; lee2024analogical) demonstrate online policy learning but require heavy reward shaping and task-specific tuning. Generative models (kim2024diffusion; hwang2025gfn) improve output diversity but often rely on statistical associations rather than interpretable abstractions. Sequential models (park2023unraveling; kim2025intention) leverage trajectory supervision to mimic human reasoning, but show stable yet limited generalization. Across paradigms, current solvers still focus on reproducing task outcomes rather than modeling reasoning processes, motivating the use of auxiliary knowledge from ARCTraj as structured cognitive supervision beyond behavioral imitation.
@@ Appendix A Examples of ARC Tasks and Reasoning
The Abstraction and Reasoning Corpus (ARC) (chollet2019ARC) evaluates abstract reasoning through grid-based transformations. Each task provides a few demonstrations from which a solver must infer an implicit rule and apply it to a test input. Solving requires recognizing structural patterns and generalizing from minimal examples without explicit instructions, serving as a benchmark for testing whether models can discover and apply abstract rules from limited evidence.
@@
Figure 7. Representative ARC tasks used in ARCTraj. Task c0f76784 (left) involves filling hollow squares with color according to their size. Task 23b5c85d (right) requires selecting the smallest rectangular object and cropping it from the grid. These tasks illustrate selection-based and compositional reasoning captured by ARCTraj trajectories. :
@@ C.1. Selection Bias in Exploration
Figure 8. Distribution of trajectory ratios across 400 ARC tasks, showing discrepancies between theoretical color requirements and actual human usage. Each subplot displays the proportion of trajectories using colors from a specific source ($\mathcal{C}^{\text{demo},\text{I}}$, $\mathcal{C}^{\text{demo},\text{O}}$, $\mathcal{C}^{\text{test},\text{I}}$, $\mathcal{C}^{\text{union}}$), with the vertical line marking how many tasks require that source. In $\mathcal{C}^{\text{demo},\text{I}}$, example input colors are often underused despite being theoretically needed, whereas $\mathcal{C}^{\text{demo},\text{O}}$ reveals frequent overuse of demonstration output colors. $\mathcal{C}^{\text{test},\text{I}}$ aligns closely with task requirements, suggesting a human bias toward test inputs. $\mathcal{C}^{\text{union}}$ confirms that all tasks are solvable using only in-task colors, consistent with ARC’s constrained design. These trends highlight inductive biases that extend beyond the availability of colors, motivating the explicit modeling of color origins. :

### v3

@@ 1. Introduction
Understanding and modeling how humans reason and solve problems, rather than reproducing only their final answers, remains a central challenge in AI (Collins et al., 2024). Such problem-solving involves conceptual abstraction, attention shifts, and the use of flexible strategies, yet these abilities remain difficult for machines to emulate (Reisberg, 2013; Ho et al., 2022). The Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) was introduced to benchmark these capabilities through grid-based tasks where solvers must infer rules from a small set of input-output examples.
@@
While the benchmark itself is well designed, most ARC research has focused on reproducing outputs rather than modeling reasoning. Early program synthesis approaches (Banburski et al., 2020; Xu et al., 2023; Ouellette, 2024; Rocha et al., 2025; Ferré, 2025) struggled to infer solution strategies because the benchmarks provide limited guidance for program construction. To address this, several studies have attempted to extract auxiliary cues from input grids or from intermediate program execution results; however, these approaches have remained task-specific and brittle. More recently, LLM-based methods (Tan and Motani, 2024; Lei et al., 2025; Pourcel et al., 2025) have leveraged prior knowledge to generate diverse candidate programs, showing improved flexibility but still facing the fundamental challenge of an enormous search space without explicit reasoning guidance.
@@
In parallel, test-time training approaches (Akyürek et al., 2025; Franzen et al., 2024; Li et al., 2025) aim to enable on-the-fly adaptation during inference; however, their learning signals still rely on static supervision rather than the evolving structure of reasoning. As a result, existing ARC approaches (e.g., program synthesis, neuro-symbolic reasoning, and test-time learning) remain limited by their reliance on static supervision and their inability to capture the unfolding of human reasoning over time.
@@
To address this limitation, we present ARCTraj, a large-scale dataset of human reasoning trajectories collected while solving ARC tasks. Each trajectory captures a temporally ordered sequence of object-level actions (i.e., moving, rotating, and flipping objects) that transform an input grid into its correct output. By capturing these multi-step trajectories, ARCTraj provides explicit reasoning supervision that connects human perception, abstraction, and strategy formation, offering a structured alternative to unconstrained program search. These logs were collected via the O2ARC web interface (Shim et al., 2024), which is designed to support natural human interaction with ARC problems. Each trajectory is annotated with metadata, including timestamps, task identifiers, and success labels, enabling both learning and analysis.
@@
Figure 1. Overview of the ARCTraj data collection process. Users solve ARC tasks through the O2ARC platform by interacting with grid-based objects. Their actions are recorded step-by-step to form semantically rich, temporally ordered trajectories.
Overview of the ARCTraj data collection process. Users solve ARC tasks through the O2ARC platform by interacting with grid-based objects. Their actions are recorded step-by-step to form semantically rich, temporally ordered trajectories.
@@
Compared to existing human ARC datasets, ARCTraj offers several advantages. Whereas H-ARC (LeGris et al., 2025) captures pixel-level edit logs and the ARC-Interactive-History-Dataset (Strandgaard, 2024) records low-level cell and operation sequences, ARCTraj provides object-centric actions with consistent formatting and semantic structure across all 400 ARC-AGI-1 training tasks. This object-level abstraction reflects how humans group meaningful visual units and apply conceptual transformations rather than pixel-level edits. Such abstraction preserves strategic intent and eliminates redundant operations, resulting in trajectories that more accurately represent human reasoning.
@@
Ultimately, ARCTraj shifts the focus of ARC research from reproducing outputs to modeling reasoning itself. By learning from human trajectories, models can acquire structured strategies that reflect System 2 thinking (Kim and Kim, 2024) and generalize across domains, such as program synthesis, robotics, and data transformation. We expect ARCTraj to promote research on transferable reasoning strategies beyond training tasks.
@@ 2. Related Work
ARC (Chollet, 2019), also known as ARC-AGI, is a benchmark designed to test human-like generalization in abstract reasoning tasks. Each task consists of a few input–output grid pairs, requiring solvers to induce and apply conceptual transformations from limited examples. It has inspired research across multiple paradigms, including program synthesis (Barke et al., 2024; Bober-Irizar and Banerjee, 2024; Butt et al., 2024), neuro-symbolic reasoning (Batorski et al., 2025; Liao and Gu, 2024; Xu et al., 2023), and test-time training (Akyürek et al., 2025; Franzen et al., 2024; Li et al., 2025), many of which were featured in the ARC Prize 2024 Technical Report (Chollet et al., 2024). However, ARC remains static; it provides only input–output pairs, offering no insight into the intermediate reasoning that drives human problem-solving. As a result, most studies evaluate solution generalization rather than the reasoning process itself. Understanding this process is crucial for modeling human-like abstraction. In this work, we use the term ARC to refer to the 400 training tasks from the ARC-AGI-1 benchmark, which form the foundation for all trajectories in ARCTraj.
@@
Several efforts have captured traces of human reasoning in ARC tasks. LARC (Acquaviva et al., 2022) collects natural language explanations, providing semantic insight but lacking action-level detail. Fast and Flexible (Johnson et al., 2021) and H-ARC (LeGris et al., 2025) record pixel-level edits, while the ARC-Interactive-History-Dataset (Strandgaard, 2024) logs cell-level actions through the BrainGridGame interface. Although these datasets represent progress, their low-level granularity and inconsistent coverage make it challenging to extract structured reasoning or integrate data into learning frameworks. They reveal what humans edit, but not how strategic reasoning evolves. A higher-level, structured view is still missing. These gaps motivate ARCTraj, which captures object-level, temporally structured trajectories aligned across all ARC tasks to support both analysis and model training.
Figure 2. Example of a single trajectory log in ARCTraj. Each action in the “actionSequence” specifies its category and operation, along with the associated grid and object state, forming a structured state–action unit.
Example of a single trajectory log in ARCTraj. Each action in the “actionSequence” specifies its category and operation, along with the associated grid and object state, forming a structured state–action unit.
@@ 3.1. Data Collection and Action Structure
Fig. 2 illustrates the trajectory structure in ARCTraj and the components of each action. Each log represents one human solving attempt and contains an ordered sequence of symbolic actions (actionSequence) recorded through the O2ARC interface (Shim et al., 2024). Every action is represented as a triplet $\langle\text{category},\text{object},\text{operation}\rangle$ associated with its grid state and timestamp.
@@
These elements define an MDP-compatible format $\langle s_{t},a_{t},s_{t+1}\rangle$, where $s_{t}$ encodes the grid configuration and $a_{t}$ represents the above triplet (Lee et al., 2024a). Formally, the reward $r_{t}\in\{1,0\}$ indicates whether the current grid $s_{t}$ matches the correct output. By preserving order and object identity, ARCTraj captures evolving reasoning processes beyond static outcomes, enabling analysis of strategies such as hypothesis testing, selective attention, and object-level abstraction.
@@ 3.2. Comparative Dataset Statistics
Among existing human ARC datasets, the most representative one is H-ARC (LeGris et al., 2025), which also records human solving processes on ARC tasks. However, the two datasets differ notably in both task coverage and the representation of actions. H-ARC collected trajectories from both the training and evaluation splits of ARC-AGI-1 and primarily logs pixel-level edits, where each action merges pixel selection and color change.
@@ 4.1. Formalizing ARC Solvers with ARCTraj
Figure 3. Overview of how ARCTraj analyses inform ARC solving. The ARC solver $f_{\theta}$ predicts $\hat{y}^{\text{test}}$ given $x^{\text{test}}$, demonstration examples $\mathcal{D}_{\text{demo}}$, and auxiliary knowledge $\mathcal{A}$ derived from human trajectories. ARCTraj provides three structured components of $\mathcal{A}$, (1) selection biases, (2) color origins, and (3) shared intentions, that capture different stages of human reasoning and can be integrated into model learning.
[…]
@@ 5. Auxiliary Knowledge from ARCTraj
To structure this investigation, we frame three research questions (RQs) that correspond to the main stages of human problem solving, from perceptual attention to hypothesis formation and strategic abstraction (Reisberg, 2013; Ho et al., 2022). Each RQ highlights a distinct aspect of reasoning behavior, together forming a coherent picture of how humans decompose, transform, and recombine patterns in ARC task solving, providing insights that bridge human cognitive processes and AI reasoning frameworks (Collins et al., 2024).
@@ 5.1. Biases in Human Grid Selections
Refer to caption Joint distribution of selection height and width.
Refer to caption Histogram of the number of selected pixels.
Figure 4. Distributions of human selection behavior in ARC tasks. Left: Selections are concentrated in compact shapes ranging from $1\times 1$ to $3\times 3$, with square- and bar-shaped regions dominating. Right: Most selections cover fewer than 20 pixels, reflecting a preference for local and perceptually salient regions.
Distributions of human selection behavior in ARC tasks. Left: Selections are concentrated in compact shapes ranging from $\times1$\$ to $\times3$\$, with square- and bar-shaped regions dominating. Right: Most selections cover fewer than 20 pixels, reflecting a preference for local and perceptually salient regions.
@@ 5.3. Shared Intentions and Strategy Patterns
Figure 5. Uniqueness analysis of human reasoning trajectories. According to the left panel, most ARC tasks show low unique trajectory ratios, indicating that human solvers often converge on similar reasoning paths. The right panel shows a representative low-uniqueness task (Task c0f76784; see Appendix A), where overlapping solution routes appear in the state-space graph.
[…]
@@
Figure 6. Preprocessing ARCTraj for downstream learning. For RL environments such as ARCLE, ARCTraj is filtered to retain only operation actions and is mapped to a Markovian state-action format using grid, object, and operation. For sequence models such as Decision Transformer, only grid and operation are used, omitting intermediate objects and environment interaction.
Preprocessing ARCTraj for downstream learning. For RL environments such as ARCLE, ARCTraj is filtered to retain only operation actions and is mapped to a Markovian state-action format using grid, object, and operation. For sequence models such as Decision Transformer, only grid and operation are used, omitting intermediate objects and environment interaction.
@@ 6.1. Current ARC Solvers and Key Trends
Recent research has explored diverse approaches to solving ARC tasks by framing visual reasoning and pattern transformation as learning or generation problems. As shown in Fig. 6, ARCTraj data are reformatted for downstream learning across solver paradigms. For reinforcement learning environments such as ARCLE (Lee et al., 2024a), trajectories are mapped to Markovian state–action pairs using grid, object, and operation. For sequence models such as Decision Transformer (Park et al., 2023), only grid and operation are retained to simplify intermediate interactions while preserving reasoning context. This unified preprocessing enables direct comparison of solvers trained on shared human reasoning traces.
@@
PPO (Lee et al., 2024a) Reinforcement Learning Solve 55–70% Demonstrated online training in ARC-like MDPs.
World Model (Lee et al., 2024b) Reinforcement Learning Solve 38–100% Enabled analogical generalization via latent dynamics.
Decision Transformer (Park et al., 2023; Kim et al., 2025) Sequential Modeling Solve 59–90% Learned trajectory-conditioned policies with inferred intentions.
Diffusion (Kim et al., 2024) Generative Modeling Solve 77–92% Generated intermediate states for plan synthesis.
GFlowNet (Hwang et al., 2024) Generative Modeling Augment 10–100% Sampled diverse goal-directed solution trajectories.
Reinforcement learning methods (Lee et al., 2024a, b) demonstrate online policy learning but require heavy reward shaping and task-specific tuning. Generative models (Kim et al., 2024; Hwang et al., 2024) improve output diversity but often rely on statistical associations rather than interpretable abstractions. Sequential models (Park et al., 2023; Kim et al., 2025) leverage trajectory supervision to mimic human reasoning, but show stable yet limited generalization. Across paradigms, current solvers still focus on reproducing task outcomes rather than modeling reasoning processes, motivating the use of auxiliary knowledge from ARCTraj as structured cognitive supervision beyond behavioral imitation.
@@ Appendix A Examples of ARC Tasks and Reasoning
The Abstraction and Reasoning Corpus (ARC) (Chollet, 2019) evaluates abstract reasoning through grid-based transformations. Each task provides a few demonstrations from which a solver must infer an implicit rule and apply it to a test input. Solving requires recognizing structural patterns and generalizing from minimal examples without explicit instructions, serving as a benchmark for testing whether models can discover and apply abstract rules from limited evidence.
@@
Figure 7. Representative ARC tasks used in ARCTraj. Task c0f76784 (left) involves filling hollow squares with color according to their size. Task 23b5c85d (right) requires selecting the smallest rectangular object and cropping it from the grid. These tasks illustrate selection-based and compositional reasoning captured by ARCTraj trajectories.
Representative ARC tasks used in ARCTraj. Task c0f76784 (left) involves filling hollow squares with color according to their size. Task 23b5c85d (right) requires selecting the smallest rectangular object and cropping it from the grid. These tasks illustrate selection-based and compositional reasoning captured by ARCTraj trajectories.
@@ C.1. Selection Bias in Exploration
Figure 8. Distribution of trajectory ratios across 400 ARC tasks, showing discrepancies between theoretical color requirements and actual human usage. Each subplot displays the proportion of trajectories using colors from a specific source ($\mathcal{C}^{\text{demo},\text{I}}$, $\mathcal{C}^{\text{demo},\text{O}}$, $\mathcal{C}^{\text{test},\text{I}}$, $\mathcal{C}^{\text{union}}$), with the vertical line marking how many tasks require that source. In $\mathcal{C}^{\text{demo},\text{I}}$, example input colors are often underused despite being theoretically needed, whereas $\mathcal{C}^{\text{demo},\text{O}}$ reveals frequent overuse of demonstration output colors. $\mathcal{C}^{\text{test},\text{I}}$ aligns closely with task requirements, suggesting a human bias toward test inputs. $\mathcal{C}^{\text{union}}$ confirms that all tasks are solvable using only in-task colors, consistent with ARC’s constrained design. These trends highlight inductive biases that extend beyond the availability of colors, motivating the explicit modeling of color origins.
[…]